OpenStudy (falling_in_katt):

What is \[3\sqrt{6}\] in radical form?

jimthompson5910 (jim_thompson5910):

It's already in radical form. Did you mean rational exponent form?

OpenStudy (falling_in_katt):

yes

jimthompson5910 (jim_thompson5910):

you can use the idea that \[\LARGE \sqrt{x} = x^{1/2}\]

jimthompson5910 (jim_thompson5910):

so that means \[\Large 3\sqrt{6} = 3*(6^{1/2})\]

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (falling_in_katt):

yes thank you

jimthompson5910 (jim_thompson5910):

you're welcome